# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ Introduction
First, we prove that $\partial u(K)$ is bounded. Since $K$ is compact in $X$ there exists $\delta>0$ such that $K^{\delta}:=\{x\in X:\ d(x,K)\leqslant\delta\}\subset X$, where $K^{\delta}$ is compact. By Lemma 1.9, $u$ is bounded on $K^{\delta}$ and by (1.7) one has
$\langle p,y-x\rangle\leqslant u(y)-u(x)\leqslant 2\max_{K^{\delta}}|u|<+\infty,\ \ \forall\,x\in K,\,y\in K^{\delta},\,p\in\partial u(x).$ (1.16)
With $y:=x+\delta p/|p|$ for $p\neq 0$ in (1.16), one has
$\frac{\delta}{|p|}\langle p,p\rangle\leqslant 2\max_{K^{\delta}}|u|,$
@@
$|p|\leqslant\frac{2}{\delta}\max_{K^{\delta}}|u|,\ \ \forall\,p\in\partial u(K),$
@@
$u(y)\geqslant u(x_{k_{j}})+\langle p_{k_{j}},y-x_{k_{j}}\rangle,\quad\forall\,y\in X,\ \forall\,j\in\mathbb{N}$
@@
where $e\vcentcolon=(y-x)/|y-x|\in\mathbb{S}^{n-1}$. Hence for $y\to x$, (1.18) shows that $u$ is differentiable in $x$ with $Du(x)=p$. This also completes the implication that $\partial u$ is single valued implies that $u$ is differentiable everywhere.
@@
The unidimensional version of Rademacher’s theorem can be proved as a corollary of the analogous result for monotone functions, which can be extended to functions of bounded variation thanks to Jordan’s decomposition, and thus to the subspace of Lipschitz continuous functions (see e.g. [41, 18]). The passage to higher dimensions essentially uses the Fubini–Tonelli theorem (see e.g. [46, Theorem 2.2.4] or [19, Theorem 3.1.6]).
@@
In addition to the result of Lemma 1.19 on sums of a convex function $u$ and a quadratic convex function $\varphi$, one has an interesting result about differences as well. It states that convex functions $u$ are differentiable at points $x$ of quadratic upper contact and is a precursor to the results on upper contact jets in Section 2 which play a key role in all that will follow.
Lemma 1.21 (Differentiability at points of upper quadratic contact).
Let $u$ be convex on $X$ and suppose there exists a quadratic function $\varphi$ such that
$(u-\varphi)(y)\leqslant 0\qquad\text{$\forall y$ near $x\in X$, with equality at $x$}.$
@@
Since $u$ is convex, $\partial u(x)$ is nonempty and hence there exists $q\in{\mathbb{R}}^{n}$ such that $u(x)+\langle q,\cdot-x\rangle\leqslant u$ on $X$. Consider now the convex function $\tilde{u}\vcentcolon=u-u(x)-\langle q,\cdot-x\rangle$ which satisfies $\tilde{u}\geqslant 0$ on $X$ and $\tilde{u}(x)=0$. Hence $0\in\partial\tilde{u}(x)$ and then, since
$\varphi=\varphi(x)+\langle D\varphi(x),\cdot-x\rangle+\tfrac{1}{2}\langle D^{2}\varphi(x)(\cdot-x),\cdot-x\rangle,\qquad\varphi(x)=u(x),$
by defining $\tilde{\varphi}\vcentcolon=\varphi-\varphi(x)-\langle q,\cdot-x\rangle$ one has
@@
and hence $\tilde{u}$ is differentiable at $x$ with $D\tilde{u}(x)=0$, which shows that $u$ is differentiable at $x$ with $Du(x)=D\varphi(x)$. Alternatively, since our argument shows that $\partial u(x)=\{D\varphi(x)\}$ and hence one knows that $u$ is differentiable at $x$ with $Du(x)=D\varphi(x)$ by Corollary 1.16. ∎
@@ 1.2. Second order theory: Alexandrov’s theorem
$\mathrm{Diff}^{2}u:=\{x\in X:\ \text{$u$ is twice differentiable in $x$}\}.$ (1.24)
@@
Given $u:X\to{\mathbb{R}}$ such that there exist $p\in{\mathbb{R}}^{n}$ and $A\in\mathcal{S}(n)$ such that (1.23) holds for $x\in X$. Then
@@
    $u$ is differentiable in $x$ with gradient $Du(x)=p$ ($p$ is unique);
@@
but choosing $y:=x+t(p-p^{\prime})$ with $t\neq 0$ so small that $y\in X$ one obtains
@@
To see that $A$ is also unique, suppose that there exist $A\neq A\in\mathcal{S}(n)$ such that (1.23) holds with $p=Du(x)$. Hence one has
@@
We conclude this first section by stating the celebrated theorem of Alexandrov on second-order differentiability of convex functions. For the convenience of the reader, a complete proof of this crucial theorem can be found in Appendix LABEL:ap:legalex, complemented by Appendix LABEL:proofsard.
@@
The proof we offer in Appendix LABEL:ap:legalex depends on two ingredients, which we can roughly describe here as a crucial property about the non-critical points of the Legendre transform of convex quadratic perturbations of a convex function (see Lemma A.7) and a Lipschitz version of Sard’s theorem (see Lemma A.9), which shows that the set of “bad” points (the critical values) has Lebesgue measure zero. For completeness, we also give a proof of Lemma A.9 in Appendix LABEL:proofsard which relies on the Besicovitch covering LemmaLABEL:app:bes, whose proof is also given. The reader is invited to have a look at Appendix LABEL:ap:legalex for a presentation of the Legendre transform and a more detailed discussion of the needed ingredients.
2. Quasi-convex functions and upper contact jets
With Alexandrov’s theorem now in hand, we will discuss the next important ingredient for the theory of viscosity solutions, which roughly speaking concerns the existence of a sufficient amount of upper (and lower) test functions. These test functions are used in the viscosity formulation of subsolutions (and supersolutions) for weakly elliptic PDEs and for the subharmonics (and superharmonics) in general potential theories. This will be systematically developped in Part II of this work, but it is perhaps useful to recall the viscosity formalism as motivation.
Given a continuous differential operator $F\in C(X\times{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n))$ on an open domain $X\subset{\mathbb{R}}^{n}$, an upper semicontinuous function $u\in\operatorname{USC}(X)$ is said to be a viscosity subsolution (on $X$) of the equation
$F[u]\vcentcolon=F(x,u(x),Du(x),D^{2}u(x))=0,\quad x\in X$ (2.1)
if and only if, for each $x\in X$ fixed, one has $F[\varphi]\geqslant 0$ for all upper test functions $\varphi$ of class $C^{2}$ for $u$ in $x$; that is, if and only if, for each $x\in X$ fixed, any $\varphi$ of class $C^{2}$ in a neighborhood of $x$ such that
$u\leqslant\varphi\quad\text{near $x$, with equality at $x$,}$ (2.2)
satisfies
$F(x,\varphi(x),D\varphi(x),D^{2}\varphi(x))\geqslant 0.$ (2.3)
Notice that an upper semicontinuous function may fail to admit an upper test function at a given point. For example, this happens for $u=|\cdot|$ at $x=0$. It is not difficult to show that there exists at least one upper test function in every neighborhood of each point of their domain. [^2] See the beginning of the proof of Proposition 3.19. However, this is not sufficient for a useful viscosity theory based on upper test functions since this guarantees the possibility to check for the validity of (2.3) only in countably many points $x$.
On the other hand, as we pointed out in the introduction, for locally quasi-convex functions one does have enough upper test functions for viscosity subsolutions by combining Alexandrov’s Theorem 2.2, the Jensen-Słodkowski Theorem 3.28 (or any of its equivalent formulations) and the Almost Everywhere Theorem 8.1. In addition, the Upper Test Jet Theorem 2.12 gives “nice continuity properties” for test jets of quasi-convex functions. Then, under certain hypotheses, the quasi-covex approximation by way of the sup-convolution allow one to pass from quasi-convex to upper semicontinuous subsolutions in viscosity theory (see LABEL:sas).
2.1. Quasi-convex functions
The definition we adopt of a (locally) quasi-convex function is the following.
@@ 2.1. Quasi-convex functions
A function $u\colon C\to{\mathbb{R}}$ is said $\lambda$-quasi-convex on the convex set $C\subset{\mathbb{R}}^{n}$ if there exists a real number $\lambda\geqslant 0$ such that the function $u+\frac{\lambda}{2}|\cdot|^{2}$ is convex on $C$.
We say that $u$ is locally quasi-convex on $X$ if for every $x\in X$ there exists a ball $B\subset X$ with $x\in B$ such that $u$ is $\lambda$-quasi-convex on $B$, for some nonnegative real number $\lambda=\lambda(x)$.
Note that we have included convexity into the definition when $\lambda=0$. In many proofs involving quasi-convex functions, we can assume without loss of generality that the functions are in fact convex. For instance, this is possible for results about differentiability, since the quadratic perturbation $\frac{\lambda}{2}|\cdot|^{2}$ is smooth. This is the case for Alexandrov’s Theorem 1.25, which clearly extends from convex to locally quasi-convex functions; that is, the following result holds.
Theorem 2.2 (Alexandrov: locally quasi-convex version).
Let $X\subset{\mathbb{R}}^{n}$ be open and let $u$ be locally quasi-convex on $X$. Then $u$ is twice differentiable (in the Peano sense) almost everywhere on $X$ with respect to the Lebesgue measure.
@@
Since the Euclidean space is Lindelöf (that is, from any open covering of a set we can extract a countable subcovering), we know that there exists a countable family of balls $\{B_{i}\}_{i\in\mathbb{N}}$ such that $u$ is $\lambda_{i}$-quasi-convex on each $B_{i}\subset X$, for some $\lambda_{i}\geqslant 0$, and $X=\bigcup_{i\in\mathbb{N}}B_{i}$. Since each $v_{i}\vcentcolon=u+\frac{\lambda_{i}}{2}|\cdot|^{2}$ is convex on $B_{i}$, by Alexandrov’s Theorem 1.25
@@
Another property that quasi-convex functions inherit from convex functions is the differentiability points of upper quadratic contact Lemma 1.21; indeed, note that if $\varphi$ is an upper contact quadratic test function for a $\lambda$-quasi-convex function $u$ at some point $x$, then $\tilde{\varphi}\vcentcolon=\varphi+\frac{\lambda}{2}|\cdot|^{2}$ is an upper contact quadratic test function for the convex function $v\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ at $x$, and one recovers the identity $Du(x)=D\varphi(x)$. This property will be used shortly in the proof of Proposition 2.5, and it will be also paraphrased in the language of upper contact jets later on in Lemma 2.13.
2.2. A quasi-convexity characterization of $C^{1,1}$
It is well-known that a function is affine if and only if it is simultaneously convex and concave. An interesting fact which helps to understand the notion of quasi-convexity is that functions which are both quasi-convex and quasi-concave must be differentiable with Lipschitz gradient (that is, of class $C^{1,1}$). Of course, we say that $u$ is quasi-concave if $-u$ is quasi-convex.
This characterization is included in [30] as an appendix, and was previously to be found in [25], in an appendix beginning as follows: “It is interesting that the condition that a function be $C^{1,1}$ is directly related to quasi-convexity, in fact it is equivalent to the function being simultaneously quasi-convex and quasi-concave. This was probably first observed by Hiriart-Urruty and Plazanet in [34]. An alternate proof appeared in [17].”
The proof exploits the fact that quasi-convexity is preserved when one convolves with a mollifier, that is a function $\eta$ such that
$\eta\in C_{\text{c}}^{\infty}({\mathbb{R}}^{n}),\quad\mathrm{supp}\,\eta\subset B_{1},\quad\eta\geqslant 0,\quad\int_{{\mathbb{R}}^{n}}\eta=1.$

### v3

@@ 1.1. First order theory
First, we prove that $\partial u(K)$ is bounded. Since $K$ is compact in $X$ there exists $\delta>0$ such that $K_{\delta}\vcentcolon=\{x\in X:\ d(x,K)\leqslant\delta\}\subset X$, where $K_{\delta}$ is compact. By Lemma 1.11, $u$ is bounded on $K_{\delta}$ and by (1.5) one has
$\langle p,y-x\rangle\leqslant u(y)-u(x)\leqslant 2\max_{K_{\delta}}|u|<+\infty,\ \ \forall\,x\in K,\,y\in K_{\delta},\,p\in\partial u(x).$ (1.12)
With $y\vcentcolon=x+\delta p/{|p|}$ for $p\neq 0$ in (1.12), one has
$\frac{\delta}{|p|}\langle p,p\rangle\leqslant 2\max_{K_{\delta}}|u|,$
@@
$|p|\leqslant\frac{2}{\delta}\max_{K_{\delta}}|u|\quad\forall\,p\in\partial u(K),$
@@
$u(y)\geqslant u(x_{k_{j}})+\langle p_{k_{j}},y-x_{k_{j}}\rangle\quad\forall\,y\in X,\ \forall\,j\in\mathbb{N}$
@@
where $e\vcentcolon=(y-x)/|y-x|\in\mathbb{S}^{n-1}$. Hence for $y\to x$, (1.14) shows that $u$ is differentiable in $x$ with $Du(x)=p$. This also completes the implication that $\partial u$ being single valued implies that $u$ is differentiable everywhere.
@@
The unidimensional version of Rademacher’s theorem can be proved as a corollary of the analogous result for monotone functions, which can be extended to functions of bounded variation thanks to Jordan’s decomposition, and thus to the subspace of Lipschitz continuous functions (see e.g. [19, 46]). The passage to higher dimensions essentially uses the Fubini–Tonelli theorem (see e.g. [20, Theorem 3.1.6] or [50, Theorem 2.2.4]).
@@
In addition to the result of Lemma 1.21 on sums of a convex function $u$ and a quadratic convex function $\varphi$, one has an interesting result about differences as well. It states that convex functions $u$ are differentiable at points $x$ of quadratic upper contact and is a precursor to the results on upper contact jets in Section 2.4 which play a key role in all that will follow.
Lemma 1.23 (Differentiability at points of quadratic upper contact).
Let $u$ be convex on $X$ (open and convex) and suppose there exists a quadratic function $\varphi$ such that
$(u-\varphi)(y)\leqslant 0\qquad\text{$\forall y$ near $x\in X$, with equality at $x$}.$ (1.19)
@@
Since $u$ is convex, $\partial u(x)$ is nonempty and hence there exists $q\in{\mathbb{R}}^{n}$ such that $u(x)+\langle q,\cdot-x\rangle\leqslant u(\cdot)$ on $X$. Consider now the convex function $\tilde{u}(\cdot)\vcentcolon=u(\cdot)-u(x)-\langle q,\cdot-x\rangle$ which satisfies $\tilde{u}\geqslant 0$ on $X$ and $\tilde{u}(x)=0$. Hence $0\in\partial\tilde{u}(x)$ and then, since
$\varphi=\varphi(x)+\langle D\varphi(x),\cdot-x\rangle+\tfrac{1}{2}\langle D^{2}\varphi(x)(\cdot-x),\cdot-x\rangle,\qquad\varphi(x)=u(x),$ (1.20)
by defining $\tilde{\varphi}(\cdot)\vcentcolon=\varphi(\cdot)-\varphi(x)-\langle q,\cdot-x\rangle$ one has
@@
and hence $\tilde{u}$ is differentiable at $x$ with $D\tilde{u}(x)=0$, which shows that $u$ is differentiable at $x$ with $Du(x)=D\varphi(x)$. Alternatively, since our argument shows that $\partial u(x)=\{D\varphi(x)\}$, one knows that $u$ is differentiable at $x$ with $Du(x)=D\varphi(x)$ by Corollary 1.18. ∎
@@
We conclude this section with some comments concerning our choice of the setting in which we consider convex functions.
Remark 1.25 (On the setting of convex functions).
We have discussed real-valued convex functions on Euclidian spaces. Much of what we have presented can be generalized to extended real-valued functions on Banach spaces. We are only interested in the finite dimensional case, where we note that Chapter 10 will show that local convexity can be defined on manifolds by using the invariance of the notion under smooth coordinate changes. It is common to consider convex functions on $X$ convex as being (,+]($\infty$,$\infty$]( - ∞ , + ∞ ]-valued functions, where +$\infty$+ ∞ is an allowed value and one keeps track of the domain ${\rm dom}(u)\vcentcolon=\{x\in X:u(x)\neq+\infty\}$. This extension has the advantage in simplifying many things, but, for us, will only have a substantial impact when we consider the Legendre transform in Appendix A. This is because the Legendre transform is more natural when defined for extended real-valued convex functions. In particular, this would avoid the pathologies that lead to the counterexample discussed in Remark A.5. Another example of the utility of considering (,+]($\infty$,$\infty$]( - ∞ , + ∞ ]-valued convex functions concerns the validity of an identity in the subdifferential of the sum of convex functions, as discussed in Lemma 1.21. Under the validity of the Attouch–Brezis condition (see [2, formula (0.3)]) the sum formula continues to hold.
On the other hand, having to keep track of the domain ${\rm dom}(u)$, where $u$ is finite, would complicate many of the statements we give. In particular, this would be true for Alexandrov’s theorem and the discussion of upper contact jets, which are essential for the viscosity theory. In addition, we want to consider locally convex functions in relation to a general potential theory, where the generalized subharmonics are [,+)[$\infty$,$\infty$)[ - ∞ , + ∞ )-valued upper semicontinuous functions. The choice of this codomain is natural for the viscosity theory for two reasons. First, one maintains the Weierstrass’s theorem: upper semicontinuous functions assume a real-valued maximum value on each compact subset. This is fundamental when using a contradiction argument for the comparison principle, amongst other things. Second, allowing the value $\infty$- ∞ gives rise to the standard viscosity theory localization of subsolutions/subharmonics by extendinng the function to be $\infty$- ∞ outside of a set of interest. This is because there are no upper contact functions and hence the extension is automaticallly a subsolution/subharmonic. We note that with this convention, an upper semicontinuous function $u$ is a viscosity subsolution of the mininal eigenvalue operator if and only if it is locally convex on the connected components where $u$ is not identically $\infty$- ∞.
@@ 1.2. Second order theory: Alexandrov’s theorem
$\mathrm{Diff}^{2}u\vcentcolon=\{x\in X:\ \text{$u$ is twice differentiable in $x$}\}.$
@@
Let $u:X\to{\mathbb{R}}$ and $x\in X$ and suppose that there exist $p\in{\mathbb{R}}^{n}$ and $A\in\mathcal{S}(n)$ such that (1.21) holds. Then the following hold:
@@
    $u$ is differentiable in $x$ with gradient $Du(x)=p$ (in particular, $p$ is unique);
@@
but choosing $y\vcentcolon=x+t(p-p^{\prime})$ with $t\neq 0$ so small that $y\in X$ one obtains
@@
To see that $A$ is also unique, suppose that there exist $A\neq A^{\prime}\in\mathcal{S}(n)$ such that (1.21) holds with $p=Du(x)$. Hence one has
@@
It is well-known that a twice differentiable locally convex function $u$ can be characterized by having positive semidefinite Hessian (see the next Proposition 1.28). In fact, this follows from the more general property that locally convex functions are those having non-negative Hessian in the viscosity sense; in other words, they are viscosity subsolutions of $\lambda_{1}(D^{2}u)=0$, where $\lambda_{1}$ denotes the minimal eigenvalue operator. This will be the content of Proposition 9.5(a), which uses the potential-theoretic formalism of Part II.
Proposition 1.28.
Let $u\colon X\to{\mathbb{R}}$ with $X\subset{\mathbb{R}}^{n}$ open. Then
[…]
and the converse is true if $\mathrm{Diff}^{2}u=X$.
Proof.
It suffices to prove the proposition under the additional assumption that $X$ is convex, so that $u$ being convex is equivalent to $u$ being locally convex (by Lemma 1.5). Indeed, if the proposition holds in that case, then it suffices to apply it to each open ball (hence convex) contained in $X$ to prove the general statement.
So, suppose that $u$ is convex in $X$ (convex). We are going to prove the more general property that, for each $x\in X$, one has $D^{2}\varphi\geqslant 0$ for each quadratic upper test function $\varphi$ for $u$ in $x$ (that is, a quadratic function satisfying (1.19)). Let $t\in[0,1)$ and let $v\in{\mathbb{R}}^{n}$ be such that $x+v\in X$. By the convexity of $u$ and the properties of $\varphi$, we have
$\begin{split}u(x+tv)\leqslant tu(x+v)+(1-t)u(x)&\leqslant t\varphi(x+v)+(1-t)\varphi(x)\\ &=\varphi(x)+t\langle D\varphi(x),v\rangle+\frac{t}{2}\langle D^{2}\varphi\,v,v\rangle,\end{split}$
where the last equality comes from the second order Taylor formula of $\varphi$ quadratic (see formula (1.20)). By Lemma 1.23 we know that $u$ is differentiable at the point $x$ of upper quadratic contact with $Du(x)=D\varphi(x)$, and $\varphi(x)=u(x)$, so the above inequalities give
$t\langle D^{2}\varphi\,v,v\rangle\geqslant u(x+tv)-u(x)-\langle Du(x),tv\rangle=o(t|v|).$ (1.23)
Since $X$ is open, we can choose any $v\in B_{\delta}(0)$ for some sufficiently small $\delta>0$, so (1.23) is equivalent to
$\langle D^{2}\varphi\,\nu,\nu\rangle\geqslant o(1)\quad\text{as}\ t\searrow 0,\quad\forall\nu\in\mathbb{S}^{n-1},$
which implies that $D^{2}\varphi$, as desired.
Now, if $u$ is twice differentiable at $x$, then in the above argument one can choose the quadratic upper test function $\varphi_{\varepsilon}(\cdot)=u(x)+\langle Du(x),\cdot-x\rangle+\frac{1}{2}\langle(D^{2}u(x)+\varepsilon I)(\cdot-x),\cdot-x\rangle$, for each $\varepsilon>0$. This yields $D^{2}u(x)\geqslant-\varepsilon I$ for all $\varepsilon>0$ and the conclusion follows by letting $\varepsilon\searrow 0$. For the converse, if $u$ is twice differentiable on $X$, then given $x,y\in X$, by the mean value theorem, there exists $z\in[x,y]$ such that
$\langle Du(x)-Du(y),x-y\rangle=\langle D^{2}u(z)(x-y),x-y\rangle\geqslant 0.$
Then it suffices to recall that the monotonicity of $Du$ characterizes the convexity of differentiable functions $u$ (Proposition 1.15). ∎
We conclude this first section by stating the celebrated theorem of Alexandrov on second-order differentiability of convex functions [1]. For the convenience of the reader, a complete proof of this crucial theorem can be found in Appendix A, complemented by Appendix B.
@@
The proof we offer in Appendix A depends on two ingredients, which we can roughly describe here as a crucial property about the non-critical points of the Legendre transform of convex quadratic perturbations of a convex function (see Lemma A.8) and a Lipschitz version of Sard’s theorem (see Lemma A.10), which shows that the set of “bad” points (the critical values) has Lebesgue measure zero. For completeness, we also give a proof of Lemma A.10 in Appendix B which relies on the Besicovitch covering Lemma B.2, whose proof is also given. The reader is invited to have a look at Appendix A for a presentation of the Legendre transform and a more detailed discussion of the needed ingredients.
Chapter 2 Semiconvex functions and upper contact jets
With Alexandrov’s theorem now in hand, this chapter will discuss the next important ingredient for the theory of viscosity solutions of differential equations and differential inclusions that involve semiconvex functions. This detailed analysis concerns upper contact jets and upper contact points of (locally) semiconvex functions and culminates in two important results: the Upper Contact Jet Theorem and the Summand Theorem. The main question concerns the existence of a sufficient amount of upper (and lower) test functions. The proof of the most delicate parts of this analysis will also make use of so-called Jensen–Słodkowski Theorem, which is the subject of the following chapter. In addition, in order to appreciate the notion of semiconvex/semiconcave functions, we will present a known characterization of Lipschitz functions with Lipschitz gradients.
2.1. Semiconvex functions
The definition we adopt of a (locally) semiconvex function is the following.
@@ 2.1. Semiconvex functions
A function $u\colon C\to{\mathbb{R}}$ is said $\lambda$-semiconvex on the convex set $C\subset{\mathbb{R}}^{n}$ if there exists a real number $\lambda\geqslant 0$ such that the function $u+\frac{\lambda}{2}|\cdot|^{2}$ is convex on $C$.
We say that $u$ is locally semiconvex on $X$ if for every $x\in X$ there exists a ball $B\subset X$ with $x\in B$ such that $u$ is $\lambda$-semiconvex on $B$, for some nonnegative real number $\lambda=\lambda(x)$.
Note that we have included convexity into the definition when $\lambda=0$. In many proofs involving semiconvex functions, we can assume without loss of generality that the functions are in fact convex. For instance, this is possible for results about differentiability, since the quadratic perturbation $\frac{\lambda}{2}|\cdot|^{2}$ is smooth. This is the case for Alexandrov’s Theorem 1.29, which clearly extends from convex to locally semiconvex functions; that is, the following result holds.
Theorem 2.2 (Alexandrov: locally semiconvex version).
Let $X\subset{\mathbb{R}}^{n}$ be open and let $u$ be locally semiconvex on $X$. Then $u$ is twice differentiable (in the Peano sense) almost everywhere on $X$ with respect to the Lebesgue measure.
@@
Since the Euclidean space is Lindelöf (that is, from any open covering of a set we can extract a countable subcovering), we know that there exists a countable family of balls $\{B_{i}\}_{i\in\mathbb{N}}$ such that $u$ is $\lambda_{i}$-semiconvex on each $B_{i}\subset X$, for some $\lambda_{i}\geqslant 0$, and $X=\bigcup_{i\in\mathbb{N}}B_{i}$. Since each $v_{i}\vcentcolon=u+\frac{\lambda_{i}}{2}|\cdot|^{2}$ is convex on $B_{i}$, by Alexandrov’s Theorem 1.29
@@
Another property that semiconvex functions inherit from convex functions is the differentiability at points of quadratic upper contact (Lemma 1.23); indeed, note that if $\varphi$ is an upper contact quadratic test function for a $\lambda$-semiconvex function $u$ at some point $x$, then $\tilde{\varphi}\vcentcolon=\varphi+\frac{\lambda}{2}|\cdot|^{2}$ is an upper contact quadratic test function for the convex function $v\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ at $x$, and one recovers the identity $Du(x)=D\varphi(x)$. This property will be used shortly in the proof of Proposition 2.5, and it will be also paraphrased in the language of upper contact jets later on in Lemma 2.13.
2.2. A semiconvexity characterization of $C^{1,1}$
It is well known that a function is affine if and only if it is simultaneously convex and concave. An interesting fact which helps to understand the notion of semiconvexity is that functions which are both semiconvex and semiconcave must be differentiable with Lipschitz gradient (that is, of class $C^{1,1}$). Of course, we say that $u$ is semiconcave if $-u$ is semiconvex.
This characterization is included in [33] as an appendix, and was previously to be found in [28], in an appendix beginning as follows: “It is interesting that the condition that a function be $C^{1,1}$ is directly related to semiconvexity, in fact it is equivalent to the function being simultaneously semiconvex and semiconcave. This was probably first observed by Hiriart-Urruty and Plazanet in [36]. An alternate proof appeared in [17].”
The proof exploits the fact that semiconvexity is preserved when one convolves with a mollifier, that is a function $\eta$ such that
$\eta\in C_{\mathrm{c}}^{\infty}({\mathbb{R}}^{n}),\quad\mathrm{supp}(\eta)\subset B_{1},\quad\eta\geqslant 0,\quad\int_{{\mathbb{R}}^{n}}\eta=1.$